# Stabilizing Reinforcement Learning with LLMs: Formulation and Practices

### v2

@@ 2.3 Token-level Objective as a First-order Approximation to Sequence-level Objective
The critical step in our formulation is to consider the following surrogate token-level optimization objective:
$\mathcal{J}^{\text{token}}(\theta)=\mathbb{E}_{x\sim\mathcal{D},\,y\sim\mu_{\theta_{\text{old}}}(\cdot|x)}\left[\sum_{t=1}^{|y|}\underbrace{\mathrm{sg}\left[\frac{\pi_{\theta}(y_{t}|x,y_{<t})}{\mu_{\theta_{\text{old}}}(y_{t}|x,y_{<t})}\right]}_{\text{token-level IS weight}}R(x,y)\,\log\pi_{\theta}(y_{t}|x,y_{<t})\right],$ (3)
which is actually the basic policy gradient algorithm (i.e., REINFORCE) equipped with a token-level IS weight. The core insight here is that we can view the token-level optimization objective in Equation (3) as a first-order approximation to the sequence-level objective in Equation (1) that we truly aim to optimize. To be specific, the gradient of this surrogate token-level objective is:
@@
Comparing Equations (2) and (4), it is obvious that $\nabla_{\theta}\,\mathcal{J}^{\text{seq}}(\theta)=\nabla_{\theta}\,\mathcal{J}^{\text{token}}(\theta)$ when $\pi_{\theta}=\mu_{\theta_{\text{old}}}$. This is why we say that Equation (3) is a first-order approximation to Equation (1). Therefore, when $\pi_{\theta}$ is close to $\mu_{\theta_{\text{old}}}$, we can improve the sequence-level objective in Equation (1) by updating the model parameters $\theta$ with the gradient in Equation (4).
@@ 4.1 MiniRL: A Minimalist Baseline Algorithm
$M_{t}=$ $\ \begin{cases}0&\text{if }\widehat{A}(x,y)>0\text{ and }r_{t}>1+\varepsilon_{\text{high}},\\ 0&\text{if }\widehat{A}(x,y)<0\text{ and }r_{t}<1-\varepsilon_{\text{low}},\\ 1&\text{otherwise},\end{cases}\quad\quad r_{t}=\frac{\pi_{\theta}(y_{t}|x,y_{<t})}{\pi_{\theta_{\text{old}}}(y_{t}|x,y_{<t})}.$
It is noteworthy that MiniRL is adopted as the baseline algorithm to maintain consistency—as closely as possible—with the token-level surrogate optimization objective in Equation 3, which has been justified by our formulation in § 2. In Appendix A, we will provide a comparison of MiniRL against other algorithms such as GRPO (Shao et al., 2024) and CISPO (Chen et al., 2025). All our experiments will be implemented uniformly based on MiniRL.
@@ 4.2 Experimental Setup
To conduct controlled experiments, we employ the standard synchronous RL framework. In each global step, we first sample a batch of $B$ prompts and sample $G$ responses for each prompt using the rollout policy in the inference engine. Then, we split the responses into NN mini-batches and apply NN gradient updates in the training engine. The finally updated policy in this global step is used as the new rollout policy in the next global step. Across all experimental runs, we use the same mini-batch size of 1,024 responses ($B=64$ and $G=16$) for each gradient update.
@@ 4.4 Results of Off-policy Training
The inference time in RL is typically bounded by the generation length and cannot be accelerated by increasing computational resources. To leverage increased compute for faster convergence, a common practice is to introduce off-policy updates. Within a synchronous RL framework, this means that a large batch of responses is split into NN mini-batches for multiple gradient updates. To investigate the recipes for stable RL training under off-policy settings, we experiment with three levels of off-policiness: with the mini-batch size fixed at 1,024 responses, the global batch size is varied to 2,048, 4,096, and 8,192, corresponding to $N=2$, 44, and 88, respectively. With MiniRL as the baseline, we compare the following methods: MiniRL (no clipping), MiniRL + R2 (no clipping), MiniRL + R2, and MiniRL + R3.

### v3

@@ 2.3 Token-level Objective as a First-order Approximation to Sequence-level Objective
The critical step in our formulation is to consider the following surrogate token-level objective:
$\mathcal{J}^{\text{token}}(\theta)=\mathbb{E}_{x\sim\mathcal{D},\,y\sim\mu_{\theta_{\text{old}}}(\cdot|x)}\left[\sum_{t=1}^{|y|}\underbrace{\frac{\pi_{\theta}(y_{t}|x,y_{<t})}{\mu_{\theta_{\text{old}}}(y_{t}|x,y_{<t})}}_{\text{token-level IS weight}}R(x,y)\right],$ (3)
with the following gradient:
@@
which is actually the basic policy gradient algorithm (i.e., REINFORCE) equipped with a token-level IS weight. The core insight here is that we can view the token-level optimization objective in Equation (3) as a first-order approximation to the sequence-level objective in Equation (1) that we truly aim to optimize. To be specific, suppose $\pi_{\theta}$ and $\mu_{\theta_{\text{old}}}$ are slightly different, let $\frac{\pi_{\theta}(y_{t}|x,y_{<t})}{\mu_{\theta_{\text{old}}}(y_{t}|x,y_{<t})}=1+\delta_{t}$ where $\delta_{t}$ is a small quantity. We have the following approximation:
$\frac{\pi_{\theta}(y|x)}{\mu_{\theta_{\text{old}}}(y|x)}=\prod_{t=1}^{|y|}\left(1+\delta_{t}\right)\approx 1+\sum_{t=1}^{|y|}\delta_{t}+O\left(\delta^{2}\right)\approx 1+\sum_{t=1}^{|y|}\delta_{t},$
where the rightmost derivation neglects second-order and higher-order small terms like $\delta_{i}\delta_{j}$. So we have:
$\nabla_{\theta}\,\mathcal{J}^{\text{seq}}(\theta)=$ $\ \mathbb{E}_{x\sim\mathcal{D},\,y\sim\mu_{\theta_{\text{old}}}(\cdot|x)}\left[R(x,y)\,\nabla_{\theta}\left(\frac{\pi_{\theta}(y|x)}{\mu_{\theta_{\text{old}}}(y|x)}\right)\right]$
$\approx$ $\ \mathbb{E}_{x\sim\mathcal{D},\,y\sim\mu_{\theta_{\text{old}}}(\cdot|x)}\left[R(x,y)\,\nabla_{\theta}\left(1+\sum_{t=1}^{|y|}\delta_{t}\right)\right]$
$=$ $\ \mathbb{E}_{x\sim\mathcal{D},\,y\sim\mu_{\theta_{\text{old}}}(\cdot|x)}\left[R(x,y)\,\nabla_{\theta}\left(\sum_{t=1}^{|y|}\frac{\pi_{\theta}(y_{t}|x,y_{<t})}{\mu_{\theta_{\text{old}}}(y_{t}|x,y_{<t})}\right)\right]$
$=$ $\ \nabla_{\theta}\,\mathcal{J}^{\text{token}}(\theta).$
This is why we say that Equation (3) is a first-order approximation to Equation (1). Therefore, when $\pi_{\theta}$ is close to $\mu_{\theta_{\text{old}}}$, we can improve the sequence-level objective in Equation (1) by updating the model parameters $\theta$ with the gradient in Equation (4).
@@ 4.1 MiniRL: A Minimalist Baseline Algorithm
$M_{t}=$ $\ \begin{cases}0&\text{if }\widehat{A}(x,y)>0\text{ and }r_{t}>1+\varepsilon_{\text{high}},\\ 0&\text{if }\widehat{A}(x,y)<0\text{ and }r_{t}<1-\varepsilon_{\text{low}},\\ 1&\text{otherwise},\end{cases}\quad\quad r_{t}=\frac{\pi_{\theta}(y_{t}|x,y_{<t})}{\pi_{\theta_{\text{old}}}(y_{t}|x,y_{<t})},$
where $\mathrm{sg}$ denotes the operation of stopping gradient. It is noteworthy that MiniRL is adopted as the baseline algorithm to maintain consistency—as closely as possible—(in gradient) with the surrogate token-level objective in Equation 3, which has been justified by our formulation in § 2. In Appendix A, we will provide a comparison of MiniRL against other algorithms such as GRPO (Shao et al., 2024) and CISPO (Chen et al., 2025). All our experiments will be implemented based on MiniRL.
@@ 4.2 Experimental Setup
To conduct controlled experiments, we employ the standard synchronous RL framework. In each global step, we first sample a batch of $B$ prompts and sample $G$ responses for each prompt using the rollout policy in the inference engine. Then, we split the responses into $N$ mini-batches and apply $N$ gradient updates in the training engine. The finally updated policy in this global step is used as the new rollout policy in the next global step. Across all experimental runs, we use the same mini-batch size of 1,024 responses ($B=64$ and $G=16$) for each gradient update.
@@ 4.4 Results of Off-policy Training
The inference time in RL is typically bounded by the generation length and cannot be accelerated by increasing computational resources. To leverage increased compute for faster convergence, a common practice is to introduce off-policy updates. Within a synchronous RL framework, this means that a large batch of responses is split into $N$ mini-batches for multiple gradient updates. To investigate the recipes for stable RL training under off-policy settings, we experiment with three levels of off-policiness: with the mini-batch size fixed at 1,024 responses, the global batch size is varied to 2,048, 4,096, and 8,192, corresponding to $N=2$, 44, and $8$, respectively. With MiniRL as the baseline, we compare the following methods: MiniRL (no clipping), MiniRL + R2 (no clipping), MiniRL + R2, and MiniRL + R3.